OpenStudy (anonymous):

3x^4+27x^2 polynomial equation

OpenStudy (anonymous):

Take out 3x^2

OpenStudy (anonymous):

i have x-0,3,-3

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

\[3x^2(x^2+9)\]

OpenStudy (anonymous):

You have to use the quadratic equation on x^2+9

OpenStudy (anonymous):

You will get an imaginary solution.

OpenStudy (anonymous):

no you dont either

OpenStudy (luigi0210):

@dabeezy remember what we did earlier with the i?

OpenStudy (anonymous):

um, i have x=0,3,-3 is this correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

do you homework, thank you but luigio is going help me

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (luigi0210):

You have to use the quad formula, unless you know how to solve without it

OpenStudy (anonymous):

dude, its not necessary

OpenStudy (anonymous):

um, i have x=0,3,-3

OpenStudy (luigi0210):

0 is right

OpenStudy (anonymous):

by factoring

OpenStudy (anonymous):

Let me just ask. Is the problem 3x^4+27x^2 or 3x^4-27x^2

OpenStudy (anonymous):

luigio, is it necessary to check answer aftr solving by factoring?

OpenStudy (anonymous):

also its immpossible to use quadratic exuation, its not in ax^2+bx+c+0 form its to the fourth exponent, not squared.....quadratic equation is used when ax^2+bx+c+0 not ax^4+bx+c+0

OpenStudy (anonymous):

no neither, its 3x^4=27x^2 solving a polynomial equation using factoring

OpenStudy (anonymous):

I was actually talking about this part. x^2+9 I got that by factoring out 3x^2 The full equation is 3x^2(x^2+9)

OpenStudy (anonymous):

move all non zero numbers to righ to obtain zero on opposing side

OpenStudy (luigi0210):

Oh I see

OpenStudy (anonymous):

You CAN use the quadratic equation on x^2+9

OpenStudy (anonymous):

its minus 9

OpenStudy (luigi0210):

Your answer is correct then, you just had us confused because of that positive

OpenStudy (anonymous):

Oh, ok.

OpenStudy (anonymous):

because it was + 27, but when u take to other side of equation it becomes negative

OpenStudy (anonymous):

but i have 3 seperate answers, is it necessart to check the solutions ?

OpenStudy (anonymous):

Tell us how you got that answer.

OpenStudy (anonymous):

in some cases will the solutions be false when plugging back in?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

It is necessary if you're not sure.

OpenStudy (anonymous):

3x^4 = 27x^2 3x^4-27x^2 = 0 3x^2(x^2-9)=0 3x^2=0 x^2-9=0 x^2=0/3 x^2=0+9 x^2=0 x^2=9 sq rt x^2=(+,-)0 sq rt x^2=sq rt 9 x=0 x=(+,-)3 x=0 x=-3 x=3

OpenStudy (anonymous):

You are right then.

OpenStudy (anonymous):

im suspicious to your answers doyourhomework

OpenStudy (anonymous):

If 3x^4 EQUALED 27x^2 in the beginning, your answers are correct.

OpenStudy (anonymous):

is 0^4=1?

OpenStudy (anonymous):

Here Plug in x = -3 \[3(-3)^2((-3)^2-9) = 0\] Plug in x = 3 \[3(3)^2(3^2-9) = 0\] Plug in x = 0 \[3(0)^2(0^2-9) = 0\]

OpenStudy (anonymous):

0^4 = 0

OpenStudy (anonymous):

0^2 = 0 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

0^23498237409234 = 0

OpenStudy (anonymous):

1^0=? 5^0=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

both?

OpenStudy (anonymous):

Yes anything to power zero = 1

OpenStudy (anonymous):

even an long retriceequation to zero power =1?

OpenStudy (anonymous):

(x^2+3x+9)^0? Like this?

OpenStudy (anonymous):

that =1?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok will u solve another wit me?

OpenStudy (anonymous):

I'll try to.

OpenStudy (anonymous):

want me repost?

OpenStudy (anonymous):

You can keep it here if you want to. Whatever you want.

OpenStudy (anonymous):

Other people can help you if you repost though

OpenStudy (anonymous):

Other than me.

OpenStudy (anonymous):

u can earn xtra medal

OpenStudy (anonymous):

It's ok. You can post here if you want.

OpenStudy (anonymous):

I don't care haha.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):